Crew 238 Operations Report 08Jan2022

[title  Operations Report – January 8th]

Good evening, here comes the

Crew 238 Operations Report 08-01-2022


Name of person filing report: Simon Werner

Non-nominal systems: none

Notes on non-nominal systems: none

Spirit rover used: No

Hours: 149.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: yes

Opportunity rover used: No

Hours: 81.6

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: No

Hours: 160.8

Beginning charge: 100%

Ending charge: 99%

Currently charging: yes

Perseverance rover used: No

Hours: unknown

Beginning charge: unknown

Ending charge: unknown

Currently charging: Yes

General notes and comments: @ Atila: can you provide me the hours and
the status of Percy? We have not used it so far, if you do not have
the hours at hand I’ll check tomorrow when I am on EVA. Checked
tunnel tarps and zip ties twice today due to windy weather, all

Summary of Hab operations:

WATER USE: 34 gallons

Water (static tank): 360

Water (loft tank): filled to 55 gallons

Water Meter: post pumping 01545623

Static to Loft Pump used – yes

Static tank pipe heater (on or off): on

Static tank heater (On or off) on

Toilet tank emptied: Yes

Summary of internet: nominal

Summary of suits and radios:.

One Suit (Nr.4) fully charged (direct via crocodile clamps) and
functional again but has no charger. The charger plug was opened as
advised by Scott. No damage to the plug was noticed. Plug reassembled,
Charger marked “do not use” and stored in the black cabinet in the EVA
Suit room.

Summary of GreenHab operations:

WATER USE: 12.5 gallons

Heater: On

Supplemental light: Yes

Harvest: Tomatoes 4.4 oz , Wheat grass 0.65 oz

Summary of Science Dome operations: Microscope (right one) used for
optical analysis of rock samples.

Dual split: Off

Summary of RAM operations: Checked battery charger plug and re-assembled it.

Summary of any observatory issues: none

Summary of health and safety issues: none

Questions, concerns and requests to Mission Support:

Is there a replacement for the broken charger of suit 4 available at
the outpost or RAM ?

Crew 238 Sol Summary 08Jan2022

[title  Sol Summary – January 8th]

Crew 238 Sol Summary Report 08Jan2022


Summary Title: Walking Mars Mindfully: Back to the Future through the Labyrinth

Author’s name: Sionade Robinson

Mission Status: Ongoing

Sol Activity Summary: Group and individual research projects
continuing. Briefing and training on drawing and walking a labyrinth –
an ancient meditative ritual for mental health and wellbeing. Session
led by crew mate Dr Kay Sandor. A temporary labyrinth marked in soft
sand north of Science Dome and completed by crew. Cleaning and
maintenance ongoing. Two walking EVAs successfully completed.

Look Ahead Plan: Visit as requested by the Mars Society expected
tomorrow between 9am and 4pm. Journalist to be hosted is Oliver
Conroy of the Guardian (UK).

Anomalies in work: None.

Weather: Sunny and warm. Cloudier than recent days and winds increasing.

Crew Physical Status: Nominal.

EVA: Morning EVA to campus area directly outside Hab to draw a
labyrinth. (Turner, Pokrywka and Sandor) in an area suggested by
Mission Support. Unfortunately the ground was too compacted.
Afternoon EVA (Robinson, Werner, Marcellino) successfully drew and
walked a temporary labyrinth in an alternative location north of
Science Dome.

Reports to be filed: Sol, Operations, two EVA Reports, Journalist and
HSO report.

Support Requested: Approval for an afternoon walking EVA (Pokrywka and
Robinson) to view and walk temporary labyrinth north of Science Dome
with journalist (see EVA Request) Please note we also plan to
initiate a crew Emergency Drill at this time (injured astronaut
rescue, previously noted in Crew Mission Plan). This will require
extraordinary approval for departure of three further team members to
recover injured astronaut from near the Science Dome back to the Hab.

Additional Support: Request collection of burnable rubbish from rear airlock.

EVA Report – January 8th

EVA # 7

Author: P.J. Marcellino

Purpose of EVA: Drawing of a labyrinth in the dirt in front of the
hab, under guidance from Master Gardener, Psychotherapist, and
certified Labyrinth Society member GHO Sandor.

Start time: 12:30:00 PM MST

End time: 01: 45:00 PM MST

Narrative: GHO Sandor conducted a mid-mission mindfulness
and wellness activity with the entire crew, starting with an attempted
drawing of a labyrinth with two crew members. Initially planned for close
to the Galileo Road intersection, we were informed by the Director that no
markings could be added to the landscape in BLM land. The Director
suggested the terrain in front of the Hab, but this turned out to be extremely
hard ground, and the drawing of the labyrinth was unsuccessful.

Destination: In front of the MDRS front airlock

Coordinates (use UTM NAD27 CONUS): 4250500 by 518250

Participants: GHO Sandor (lead), HSO Turner, ART Pokrywka

Road(s) and routes per MDRS Map: None

Mode of travel: By foot

Copyright © The Mars Society. All rights reserved. | Main Site